# SlowTrack: Increasing the Latency of Camera-based Perception in Autonomous Driving Using Adversarial Examples

### 15 Dec 2023 v1

@@ 1 Initialization: $\mathcal{T}\leftarrow\emptyset$ for frame $f_{k}$ in V do
       $\mathcal{D}_{k}\leftarrow\texttt{Det}(f_{k})$downarrow $O(n^{2})$ $or$ $O(n)$ $\mathcal{D}_{reserved}\leftarrow{F}({D}_{k})$downarrow $O(m)$ /* predict locations of trackers */
@@ 30             $\mathcal{T}_{unconfirmed}\leftarrow\mathcal{T}_{unconfirmed}\cup\{d\}$
Formulation. Detectors often restrict the maximum number of detection boxes. This constraint results in a worst-case time complexity for the detection module of $O(\mathcal{|D|}_{max}^{2})$ , effectively reducing the impact of previous latency attacks on object tracking (Chen et al. 2023; Shapira et al. 2023) Thus, we propose an attack methodology that focuses on increasing the latency of the subsequent tracking stage under the constraint of limiting the maximum number of detection boxes. For a camera-based perception pipeline $P$, given the original image $x$, the attack goal is to craft an adversarial example $x^{*}$ to maximize the tracking pipeline latency, while keeping the added adversarial perturbations imperceptible. We formulate it as the following optimization problem:
@@
, where $T$ indicates the time function. Since the worst-case time complexity of the tracking stage is $O(\mathcal{|D|}_{max}^{2})$, we can only increase the latency by increasing the number of trackers $m$. Our attack strategy is designed to create detection boxes that exploit vulnerabilities in the tracker management module of AD. Specifically, for avoiding the temporary loss of objects in consecutive video frames due to occlusion, etc., the lost tracker will not be deleted immediately until the object is lost for $R$ consecutive frames. Leveraging this mechanism, we are able to inject more and more trackers into the tracking module by strategically creating detection boxes through carefully designed perturbations. The detection boxes that appear in each frame are not associated with existing trackers and cause new tracking boxes to be created. As a result, the worst-case time complexity of the tracking module under our attack method is $O(R\mathcal{|D|}_{max}^{2})$.
@@
The overview of SlowTrack is in Fig. 1. Assuming that we divide the image into 3 regions, in frame 0, we select the top region and make the detection boxes appear in it, which are initialized as new track boxes. Similarly, in frame $H=1$ we make the fake detection boxes appear in the middle region, not associated with the existing track boxes, and make the tracker management module create new track boxes, while the track boxes in frame 0 do not disappear. These frames are not associated with existing frames, making the tracker management module create new trackers, and the trackers initialized in frame 0 are not deleted, allowing to inject $\mathcal{|D|}_{max}$ trackers into the track__\__pool. Meanwhile, we make the detection boxes of the corresponding region reappear before the $R$ frames to keep the trackers not deleted. Since the tracker management module only initializes the detection boxes that detected $H$ frames consecutively, our attack strategy can be formulated as a scheduling problem: divide the image into $K$ regions, each of which needs to be selected for $H$ frames consecutively in the initialization stage, and then be selected once more before $R$ frames. We use the greedy algorithm with the maximum value of $K$: $(R-(H-2))$.
@@ Acknowledgments
We would like to thank Ziwen Wan, Yunpeng Luo, Tong Wu, Xinyang Zhang, and the anonymous reviewers for their valuable and insightful feedback. This research was supported in part by National Key R&D Program of China (2020AAA0107702); the NSF under grants CNS-1929771, CNS-2145493, and CNS-1932464; National Natural Science Foundation of China (U21B2018, 62161160337, 62132011, 62376210, 62006181, U20B2049); Shaanxi Province Key Industry Innovation Program (2021ZDLGY01-02); and Fundamental Research Funds for the Central Universities under grant (xtr052023004, xtr022019002). Chao Shen is the corresponding author.

### 26 Dec 2023 v2

@@ 1 Initialization: $\mathcal{T}\leftarrow\emptyset$ for frame $f_{k}$ in V do
       $\mathcal{D}_{k}\leftarrow\texttt{Det}(f_{k})$downarrow $O(n^{2})$ + $O(n)$ $\mathcal{D}_{reserved}\leftarrow{F}({D}_{k})$downarrow $O(m)$ /* predict locations of trackers */
@@ 30             $\mathcal{T}_{unconfirmed}\leftarrow\mathcal{T}_{unconfirmed}\cup\{d\}$
Formulation. Detectors often restrict the maximum number of detection boxes. This constraint results in a worst-case time complexity for the detection module of $O(\mathcal{|D|}_{max}^{2})$ , effectively reducing the impact of previous latency attacks on object tracking (Chen et al. 2023; Shapira et al. 2023). Thus, we propose an attack methodology that focuses on increasing the latency of the subsequent tracking stage under the constraint of limiting the maximum number of detection boxes. For a camera-based perception pipeline $P$, given the original image $x$, the attack goal is to craft an adversarial example $x^{*}$ to maximize the tracking pipeline latency, while keeping the added adversarial perturbations imperceptible. We formulate it as the following optimization problem:
@@
, where $T$ indicates the time function. Since the worst-case time complexity of the detection stage is $O(\mathcal{|D|}_{max}^{2})$, we can only increase the latency by increasing the number of trackers $m$. Our attack strategy is designed to create detection boxes that exploit vulnerabilities in the tracker management module of AD. Specifically, for avoiding the temporary loss of objects in consecutive video frames due to occlusion, etc., the lost tracker will not be deleted immediately until the object is lost for $R$ consecutive frames. Leveraging this mechanism, we are able to inject more and more trackers into the tracking module by strategically creating detection boxes through carefully designed perturbations. The detection boxes that appear in each frame are not associated with existing trackers and cause new tracking boxes to be created. As a result, the worst-case time complexity of the tracking module under our attack method is $O(R\mathcal{|D|}_{max}^{2})$, where $m=R\mathcal{|D|}_{max}$ and $n^{\prime}=|D|_{max}$. Thus, it can lead to more computation cost than detection attacks in prior works..
@@
The overview of SlowTrack is in Fig. 1. Assuming that we divide the image into 3 regions, in frame 0, we select the top region and make the detection boxes appear in it, which are initialized as new track boxes. Similarly, in frame $H=1$ we make the fake detection boxes appear in the middle region, not associated with the existing track boxes, and make the tracker management module create new track boxes, while the track boxes in frame 0 do not disappear. These frames are not associated with existing frames, making the tracker management module create new trackers, and the trackers initialized in frame 0 are not deleted, allowing to inject $\mathcal{|D|}_{max}$ trackers into the track__\__pool. Meanwhile, we make the detection boxes of the corresponding region reappear before the $R$ frames to keep the trackers not deleted. Since the tracker management module only initializes the detection boxes that detected $H$ frames consecutively, our attack strategy can be formulated as a scheduling problem: divide the image into $K$ regions, each of which needs to be selected for $H$ frames consecutively in the initialization stage, and then be selected once more before $R$ frames. We use the greedy algorithm with the maximum value of $K$: $R-H+1$ (when $R-K=H-1$, no region can be added).
@@ Acknowledgments
We would like to thank Ziwen Wan, Yunpeng Luo, Tong Wu, Xinyang Zhang, and the anonymous reviewers for their valuable and insightful feedback. This research was supported in part by National Key R&D Program of China (2020AAA0107702); the NSF under grants CNS-1929771, CNS-2145493, and CNS-1932464; National Natural Science Foundation of China (U21B2018, 62161160337, 62132011, 62376210, 62006181, U20B2049); Shaanxi Province Key Industry Innovation Program (2021ZDLGY01-02); and Fundamental Research Funds for the Central Universities under grant (xtr052023004, xtr022019002). Chao Shen and Qi Alfred Chen are the corresponding authors.